OpenStudy (anonymous):

z = 2/3 for z - 1/5 = 5/8 True or false t = 9 for 10 ½ = t + 1 ½ True or false

OpenStudy (er.mohd.amir):

z-1/5=2/3-1/5 take LCM and solve it 10 and 1/2=21/2 1 and 1/2=3/2 so t=21/2-3/2